lynnparker:

Evaluate. [100 ÷ (27 – 22)] – 3 • 5 + 25

finny:

[100:(27-22)]-3·5+25= (100:5)-15+25= 20-15+25= 5+25=? will give you your answer

lynnparker:

thx

finny:

yw